OpenStudy (anonymous):

How can I find g(x) in this equation using the given information? Usually I'm better at this kind of stuff, but this one really stumped me. h(x)=(f o g)(x) <--- I know what that means And I'll put the rest of it in the comments because I can't make the right equation here in the question

OpenStudy (anonymous):

\[h(x)=^{4}\sqrt{x+7}\]and\[f(x)=^{4}\sqrt{x+1}\]

OpenStudy (anonymous):

I don't know how to find g(x)...

OpenStudy (phi):

h(x)=(f o g)(x) means h(x)= f ( g(x) ) in other words, wherever you see an x in f(x), replace it with g(x)

OpenStudy (phi):

\[\sqrt[4]{x+7}= \sqrt[4]{g(x)+1}\]

OpenStudy (anonymous):

Yes, I know that much but I need to find g(x).

OpenStudy (anonymous):

Where do I go from there? lol this is by far one of the most confusing questions I've gotten this year.

OpenStudy (phi):

it must be that x+7 = g(x)+1 solve for g(x)

OpenStudy (anonymous):

Ohh okay, so then x+6 = g(x), because I moved the one to the other side by subtraction?

OpenStudy (phi):

yes. as a check, with g(x)= x+6 what is (f o g)(x) ?

OpenStudy (anonymous):

Okay, since (f o g)(x) = f(g(x)), then it would be\[^{4}\sqrt{(x+6)+1}\]

OpenStudy (phi):

which matches h(x)

OpenStudy (phi):

if you simplify x+6+1 to x+7

OpenStudy (anonymous):

Thank you very much (:

OpenStudy (phi):

yw